A manufacturing plant produces fixed size square and circular metal panels.
The mass of a square panel is 13kg and that of a circular panel is 22 kg.
Use algebraic techniques to find how many of each panel there are in a truck
loaded with 65 panels of total mass 1205 kg.
I know the answers are 13kg=25 and 22kg=40
but i am struggling with how to use algebraic techniques, please help me understand?!?!

Answer by rothauserc(4718)

Let x be the number of 13kg panels and y be the number of 22kg panels, then
x + y = 65
(x*13) + (y*22) = 1205
solve for x in first equation
x = 65 - y
(65-y)*13 + (y*22) = 1205
845 -13y + 22y = 1205
9y = 360
y = 40 and
x = 25
therefore
there are 25 13kg panels and 40 22kg panels